OpenStudy (anonymous):

solve the system of equations!! 3x+4y=10 -6x-8y+-20

OpenStudy (anonymous):

it is an equivalent equation. There are infinite solutions

OpenStudy (anonymous):

x = (10-4y)/3 y = (10 - 3x)/4

OpenStudy (anonymous):

There is no numerical solution

OpenStudy (anonymous):

You can use substitution or addition to solve these equations or use row reduction